The height of a person is 1.65 m. Express it into cm and mm.


Grade:upto college level

6 Answers

Jitender Pal
10 years ago

Height of the person = 1.65 m

1 m = 100 cm

1.65 m = 100 × 1.65 = 165 cm

Hence, the height of the person is 165 cm.

Again, 1 m = 100 cm = 1000 mm

Therefore, 1.65 m = 1.65 × 10 = 1650 mm

Hence, the height of the person is 1650 mm.

Adarsh
8 years ago

Height of the person = 1.65 m

1.65 m = 100 × 1.65 = 165 cm
1.65 m = 1.65 × 10 = 1650 mm
Vasanth Bhaskara
8 years ago
1.65 m=100×1.65 =165 cm
1.65 m=1.65×10 =1650 mm
Ayan
6 years ago
Given height of a person = 1.65
we know that 1 m = 100 cm , 1 m = 1000 mm
1.65*100 = 165, 1.65*1000 = 1650 cm
praveen kumar
6 years ago
given height=1.65 m
we have 1m = 100 cm
1.65 m=165 cm
and then 1 cm =10 mm
165 cm=1650 mm
1.65 m=165 cm=1650 mm
 
ankit singh
3 years ago
 

1 m = 100 cm

1.65 m = 100 × 1.65 = 165 cm

Hence, the height of the person is 165 cm.

Again, 1 m = 100 cm = 1000 mm

Therefore, 1.65 m = 1.65 × 10 = 1650 mm